JET-COOLED, BROAD RANGE NEAR-IR SCAN OF REACTIVE INTERMEDIATES USING CAVITY RINGDOWN SPECTROSCOPY

\underline{NEAL D. KLINE}, TERRANCE J. CODD, MING-WEI CHEN^{a}, And TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210..

A technique has been developed for obtaining broad range scans of jet-cooled radicals using cavity ringdown spectroscopy. The talk will describe a method that uses a slit jet expansion in tandem with an electrical discharge to produce the reactive intermediates and obtain rotational temperatures of 15-30 K and effective vibrational temperatures of 0 K. The spectrum can be recorded by scanning the second stokes of an H_{2} Raman shifted YAG-pumped dye laser with bandwidth of ≈ 0.1 cm^{−1}. This technique has been used to obtain the jet-cooled Ã− X̃ spectra of the NO_{3} radical and the 2,1-hydroxypropyl peroxy radical (2,1-HPP). Obtaining the jet-cooled spectra helped to identify hot bands present in the room temperature spectrum of NO_{3} and also helped to identify cold, low-frequency fundamental bands present in the 2,1-HPP spectrum.

aPresent Address: Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL, 61801